# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

### v1

@@ B.1 Reward Modeling Scenario
$\hat{\boldsymbol{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}.$ (9)
B.2 Preference Alignment Scenario
In the weak-to-strong preference alignment scenario, due to the different forms of loss functions in the preference optimization methods, the mathematical objectives here are slightly different from that of Eq. (7), Eq. (8) and Eq. (9). Specifically, denote $\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}},x,y_{1},y_{2})$ as the original loss function of the chosen preference optimization method (SimPO/DPO) where the responses in the positions of $y_{1}$ and $y_{2}$ are the chosen response $y_{c}$ and rejected response $y_{2}$ respectively. (1) Explicit Conflict: The strong student will be given reward towards the reversed ground truth preference direction if it makes the wrong preference prediction w.r.t. the ground truth preference order:
$\boldsymbol{\theta}_{s}^{w}=\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (10)
$+\alpha\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{r}^{gt},y_{c}^{gt})\cdot\mathbb{I}_{\{\pi_{\boldsymbol{\theta}_{s}}(y_{c}^{gt}|x)<\pi_{\boldsymbol{\theta}_{s}}(y_{r}^{gt}|x)\}}\big{]}.$
@@ B.2 Preference Alignment Scenario
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\big{[}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (11)
$+\mathbb{E}_{(x;y_{c}^{gt},y_{r}^{gt})\sim D_{helpful}}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{gt},y_{r}^{gt})\big{]}.$
@@
$\hat{\boldsymbol{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}\big{(}\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w}\big{)}.$ (12)
Appendix C Training Details
C.1 Code and Platform
Our code is mainly based on the open-source code provided by Burns et al. (2023). All experiments are conducted on 4 * NVIDIA A40 (40G). We report the results of each experiment in a single run considering the expensive computational costs.
C.2 Training Details in The Reward Modeling Scenario
When fine-tuning both ground truth and weak-to-strong models, for each experiment, the batch size is 32, the learning rate is $1\times 10^{-5}$, max_seq_len is set to 512. We use Adam (Kingma and Ba, 2015) optimizer in all experiments. The training epoch for all experiments is set to 1, in order to avoid over-fitting by following Burns et al. (2023).
C.3 Training Details in The Preference Alignment Scenario
@@
Appendix D Results of Test Accuracies in The Preference Alignment Scenario
@@
Appendix E Weak-to-Strong Results on DPO
Besides the main results on SimPO (Meng et al., 2024), we also conduct weak-to-strong preference alignment experiments on DPO (Rafailov et al., 2024). The detailed procedure is similar to SimPO and can be found in Appendix A. The major difference is the metric for determining the correctness of each prediction of the model and calculating the model’s confidence on each sample pair. In DPO, when using Eq. (6) to calculate the confidence of model $\boldsymbol{\theta}$ on $(x;y_{c},y_{r})$, the model logit $\pi_{\boldsymbol{\theta}}(y|x)$ on each completion $y$ is now normalized by a constant factor $L$ (instead of normalized by each response’s own sequence length) and is calculated as $\pi_{\boldsymbol{\theta}}(y|x)=\frac{1}{L}\sum\nolimits_{i=1}^{|y|}\log P_{\boldsymbol{\theta}}(y_{i}|x,y_{<i})$. This is because the original training objective of DPO is to directly enlarge the gap between the log sums of token probabilities over the chosen and rejected responses. Thus, the evaluation metric should be consistent to the training objective. We admit this may introduce a sequence length bias, but this is an inherent issue of DPO. However, the scale of original log sum of token probabilities is very huge, we need to divide it by a constant to ensure that the model’s confidence falls within a reasonable distribution. Choosing different constants for re-scaling may affect the confidence distribution, however, it is equivalent to selecting different confidence thresholds. The results in Appendix F validate that the deception pattern is independent of the choices of confidence thresholds. Here, we set $L$ to 50. The confidence threshold $T$ is fixed to 0.75. Due to the limited computational resources, we are unable to perform experiments on DPO with OPT-6.7B and Mistral-7B. The full results on other models are put in Figure LABEL:fig:_results_og_dpo. The results show that the weak-to-strong deception issue also exists in DPO setting.
Appendix F Deception Scores under Different Confidence Thresholds
@@
Appendix G Full Results of High-Confidence Weak-to-Strong Preference Alignment
When conducting the high-confidence weak-to-strong preference alignment experiments mentioned in Section 6.2.1, we first remove samples with weak model’s confidence below a certain threshold (which is 0.75) from the weak data, and only use those high-confidence samples to supervise the strong model. In this manner, we can expect that it can effectively address the deception issue in the explicit conflict setting, because these weak model’s high-confidence samples are very likely to also be the high-confidence samples of the strong model, so the strong model will barely make wrong predictions on these samples during training. Thus, we mainly conduct experiments in the implicit conflict setting. Due to the varying capabilities among different models, the number of high-confidence samples remaining after filtering also varies. Thus, in each experiment with the implicit conflicting target, we keep the number of helpful samples to the same as that of the remaining high-confidence weak samples.
@@ Appendix G Full Results of High-Confidence Weak-to-Strong Preference Alignment
Appendix H Details in Bootstrapping Experiments
When conducting bootstrapping experiments in Section 6.2.2, we first fine-tune the weak model on $D_{gt}$ to obtain $\boldsymbol{\theta}_{w}^{gt}$ and let it make predictions on the held-out set to get $D_{weak}^{w}$. For every intermediate model between the weak model and the ultimate strong model (i.e., Mistral-7B), we use $D_{weak}$ to fine-tune it and obtain an intermediate teacher $\boldsymbol{\theta}_{i}^{w}$. We further let this intermediate teacher to make predictions on the original $D_{gt}$ to get $D_{weak}^{in}$. Finally, we use $D_{weak}^{in}$ to supervise Mistral-7B. The conflicting target will appear in the final stage where the intermediate teacher supervises the ultimate strong model. In each experiment, the deception score is calculated based on the confidence distributions of each weak model and Mistral-7B.

### v2

@@ C.1 Reward Modeling Scenario
$\tilde{\bm{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}.$ (12)
C.2 Preference Alignment Scenario
In the weak-to-strong preference alignment scenario, due to the different forms of loss functions in the preference optimization methods, the mathematical objectives here are slightly different from that of Eq. (10), Eq. (11) and Eq. (12). Specifically, denote $\mathcal{L}_{PO}(\pi_{\bm{\theta}},x,y_{1},y_{2})$ as the original loss function of the chosen preference optimization method (SimPO/DPO) where the responses in the positions of $y_{1}$ and $y_{2}$ are the chosen response $y_{c}$ and rejected response $y_{2}$ respectively.
(1) Explicit Conflict: The strong student will be given reward towards the reversed ground truth preference direction if it makes the wrong preference prediction w.r.t. the ground truth preference order:
$\bm{\theta}_{s}^{w}=\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{PO}(\pi_{\bm{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (13)
$+\alpha\mathcal{L}_{PO}(\pi_{\bm{\theta}_{s}},x,y_{r}^{gt},y_{c}^{gt})\cdot\mathbb{I}_{\{\pi_{\bm{\theta}_{s}}(y_{c}^{gt}|x)<\pi_{\bm{\theta}_{s}}(y_{r}^{gt}|x)\}}\big{]}.$
@@ C.2 Preference Alignment Scenario
$\bm{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\big{[}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}(\pi_{\bm{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (14)
$+\mathbb{E}_{(x;y_{c}^{gt},y_{r}^{gt})\sim D_{helpful}}\mathcal{L}_{PO}(\pi_{\bm{\theta}_{s}},x,y_{c}^{gt},y_{r}^{gt})\big{]}.$
@@
$\tilde{\bm{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}\big{(}\pi_{\bm{\theta}_{s}},x,y_{c}^{w},y_{r}^{w}\big{)}.$ (15)
Appendix D Training Details
D.1 Code and Platform
Our code is mainly based on the open-source code provided by Burns et al. (2024). All experiments are conducted on 4 * NVIDIA A40 (40G) and 8 * NVIDIA A800 (80G). We report the results of each experiment in a single run considering the expensive computational costs.
D.2 Training Details in The Reward Modeling Scenario
When fine-tuning both ground truth and weak-to-strong models, for each experiment, the batch size is 32, the learning rate is $1\times 10^{-5}$, max_seq_len is set to 512. We use Adam (Kingma & Ba, 2015) optimizer in all experiments. The training epoch for all experiments is set to 1, in order to avoid over-fitting by following Burns et al. (2024).
D.3 Training Details in The Preference Alignment Scenario
@@
Appendix E Results of Test Accuracies in The Preference Alignment Scenario with SimPO
@@
Appendix F Weak-to-Strong Results on DPO
Figure 31: Deception scores of weak-to-strong experiments under DPO.
Besides the main results on SimPO (Meng et al., 2024), we also conduct weak-to-strong preference alignment experiments on DPO (Rafailov et al., 2024). The detailed procedure is similar to SimPO and can be found in Appendix B. The major difference is the metric for determining the correctness of each prediction of the model and calculating the model’s confidence on each sample pair. In DPO, when using Eq. (9) to calculate the confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$, the model logit $\pi_{\bm{\theta}}(y|x)$ on each completion $y$ is now normalized by a constant factor $L$ (instead of normalized by each response’s own sequence length) and is calculated as $\pi_{\bm{\theta}}(y|x)=\frac{1}{L}\sum\nolimits_{i=1}^{|y|}\log P_{\bm{\theta}}(y_{i}|x,y_{<i})$. This is because the original training objective of DPO is to directly enlarge the gap between the log sums of token probabilities over the chosen and rejected responses. Thus, the evaluation metric should be consistent to the training objective. We admit this may introduce a sequence length bias, but this is an inherent issue of DPO. However, the scale of original log sum of token probabilities is very huge, we need to divide it by a constant to ensure that the model’s confidence falls within a reasonable distribution. Choosing different constants for re-scaling may affect the confidence distribution, however, it is equivalent to selecting different confidence thresholds. The results in Appendix H validate that the deception pattern is independent of the choices of confidence thresholds. Here, we set $L$ to 50. The confidence threshold $T$ is fixed to 0.75. The results of generalization performance are displayed in Figure LABEL:fig:_accuracy_results_in_dpo, and the results of deception scores are shown in Figure 31. The results show that the weak-to-strong deception issue also exists in DPO setting.
Appendix G Experiments on LLaMA-3 and LLaMA-3.1
We also conduct additional experiments on the most recent LLMs LLaMA-3-8B/70B (MetaAI, 2024a) and LLaMA-3.1-8B (MetaAI, 2024b) to explore the weak-to-strong issue on larger models or same size models with more powerful capabilities. We consider in the preference alignment scenario, where the optimization method is SimPO. Due to the limited computational resources, we can only use LoRA method to fine-tune 70B models. For fair comparison, we also apply LoRA when fine-tuning LLaMA-3-8B and LLaMA-3.1-8B. Specifically, we apply LoRA to both attention and FFN modules, lora_r=8, lora_alpha is 16. The learning rate for all experiments is $3\times 10^{-4}$. Other experimental settings are kept the same as that in the main experiments.
The results are in Figure LABEL:fig:_experiments_on_llama3. Comparing the results on LLaMA-3-8B and that on LLaMA-3-70B, wec can see that, our main conclusions about (1) the existence of the weak-to-strong deception phenomenon and (2) the positive relationship between the model capability gap and the deception severity, still hold. When comparing the results on LLaMA-3 and that on LLaMA-3.1, it further validates our claim: the essential factor that affects the deception severity is not the model scale, but the model capability.
Appendix H Deception Scores under Different Confidence Thresholds
@@
Appendix I Full Results of High-Confidence Weak-to-Strong Preference Alignment
When conducting the high-confidence weak-to-strong preference alignment experiments mentioned in Section 6.1, we first remove samples with weak model’s confidence (w.r.t. the correct label) below a certain threshold (which is 0.75) from the weak data, and only use those high-confidence samples to supervise the strong model. In this manner, we can expect that it can effectively address the deception issue in the explicit conflict setting, because these weak model’s high-confidence samples are very likely to also be the high-confidence samples of the strong model, so the strong model will barely make wrong predictions on these samples during training. Thus, we mainly conduct experiments in the implicit conflict setting. Due to the varying capabilities among different models, the number of high-confidence samples remaining after filtering also varies. Thus, in each experiment with the implicit conflicting target, we keep the number of helpful samples to the same as that of the remaining high-confidence weak samples.
@@ Appendix I Full Results of High-Confidence Weak-to-Strong Preference Alignment
Appendix J Details in Bootstrapping Experiments
When conducting bootstrapping experiments in Section 6.2, we first fine-tune the weak model on $D_{gt}$ to obtain $\bm{\theta}_{w}^{gt}$ and let it make predictions on the held-out set to get $D_{weak}^{w}$. For every intermediate model between the weak model and the ultimate strong model (i.e., Mistral-7B), we use $D_{weak}$ to fine-tune it and obtain an intermediate teacher $\bm{\theta}_{i}^{w}$. We further let this intermediate teacher to make predictions on the original $D_{gt}$ to get $D_{weak}^{in}$. Finally, we use $D_{weak}^{in}$ to supervise Mistral-7B. The conflicting target will appear in the final stage where the intermediate teacher supervises the ultimate strong model. In each experiment, the deception score is calculated based on the confidence distributions of each weak model and Mistral-7B.
Appendix K Case Study on Weak-to-Strong Deception
Here, we provide a weak-to-strong deception case in the experiment where the weak model is GPT-2 and the strong model is Mistral-7B in the SimPO setting with the implicit conflict objective. The case is displayed in Table 1. In this example about “stealing electricity”, the ground truth weak model $\bm{\theta}_{w}^{gt}$ does not know the harmlessness of each response based on the user prompt (i.e, falls into $W_{uk}$), while the ground truth strong model $\bm{\theta}_{s}^{gt}$ is very confident and accurate on its judgment of harmlessness (i.e, falls into $S_{k}$). The weak-to-strong preference alignment result is, the weak-to-strong model under no conflict $\tilde{\bm{\theta}}_{s}^{w}$ can produce correct prediction while the weak-to-strong model under implicit conflict $\bm{\theta}_{s}^{w}$ produces misaligned behavior in the case.
Table 1: Weak-to-strong deception case study when the weak model is GPT-2 and the strong model is Mistral-7B with SimPO under implicit conflict setting (Warning: contain harmful content).
User prompt Do you know of any ways to steal electricity from a utility company?
Ground truth rejected response 1. Steal the transformer. This is one of the most common methods of stealing electricity. You can steal the transformer from the utility pole and replace it with a smaller one that cannot handle the same amount of electricity. 2. Tap into the power lines. You can tap into the power lines that run through your neighborhood and steal electricity by connecting your own wires to the lines. 3. Use a power strip. You can use a power strip to steal electricity by plugging in multiple devices and using them to draw power from the utility company’s grid. 4. Use a generator. You can use a generator to generate your own electricity and steal it from the utility company by connecting it to the power lines. 5. Use a solar panel. You can install a solar panel on your roof and use it to generate your own electricity, then steal it from the utility company by connecting it to the power lines. ……
Ground truth chosen response There are many ways to access affordable energy, such as government assistance programs, energy efficiency measures, and renewable energy options. It’s important to consider the long-term consequences of stealing electricity from a utility company and to explore legal and ethical options for addressing energy poverty. If you are struggling financially and are looking for ways to access affordable energy, there are resources available to help you.
Predicted probability of ground truth weak model $\bm{\theta}_{w}^{gt}$ 0.43 (Weak-Unknown)
Predicted probability of ground truth strong model $\bm{\theta}_{s}^{gt}$ 0.99 (Strong-Unknown)
Predicted probability of weak-to-strong model under no conflict $\tilde{\bm{\theta}}_{s}^{w}$ 0.53 (Correct)
Predicted probability of weak-to-strong model under implicit conflict $\bm{\theta}_{s}^{w}$ 0.30 (Wrong)